# MOCHA: Multi-modal Objects-aware Cross-arcHitecture Alignment

### v3

@@ Abstract
We introduce MOCHA (Multi-modal Objects-aware Cross-arcHitecture Alignment), a knowledge distillation approach that transfers region-level multimodal semantics from a large vision-language teacher (e.g., LLaVa) into a lightweight vision-only object detector student (e.g., YOLO). A translation module maps student features into a joint space, where the training of the student and translator is guided by a dual-objective loss that enforces both local alignment and global relational consistency. Unlike prior approaches focused on dense or global alignment, MOCHA operates at the object level, enabling efficient transfer of semantics without modifying the teacher or requiring textual input at inference. We validate our method across four personalized detection benchmarks under few-shot regimes. Results show consistent gains over baselines, with a +10.1 average score improvement. Despite its compact architecture, MOCHA reaches performance on par with larger multimodal models, proving its suitability for real-world deployment.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) such as CLIP (Radford et al. 2021), Flamingo (Alayrac et al. 2022), and LLaVa (Liu et al. 2023a) have demonstrated remarkable zero-shot and open-vocabulary capabilities, owing to their rich representations and scale. These models, however, are typically large and computationally intensive, limiting their applicability in real-time or resource-constrained scenarios such as mobile devices (e.g., smartphones or robots). In contrast, lightweight architectures such as YOLO (Jocher, Chaurasia, and Qiu 2023) offer fast and memory-efficient object detection, at the cost of degraded performance in low-data regimes and vulnerability to neural collapse (Kothapalli 2022; Papyan, Han, and Donoho 2020), where features become indistinguishably aligned, losing their discriminative power. In this work, we aim to bridge this gap by proposing MOCHA, a knowledge distillation approach that transfers object-centric multimodal embeddings from a vision-language teacher (e.g., LLaVa) into a compact vision-only student detector. MOCHA exploits the property that semantically related concepts exhibit similar structures in the embedding space across different modalities under a well-generalized backbone (Huh et al. 2024; Girdhar et al. 2023; Camuffo, Michieli, and Milani 2023). By aligning and regularizing the student’s feature space, it shows improved generalization in few-shot detection settings (Fig. 1). Our approach is built on three stages that include (1) pretraining of a student model, (2) knowledge distillation on rich joint visual and textual features from a frozen teacher, and (3) few-shot personalization with frozen student model and prototype-based classifier (Snell, Swersky, and Zemel 2017).
@@
Our contribution lies primarily in the knowledge distillation stage, where we introduce: (i) extraction of rich joint visual-language representations conditioned on image regions and class labels from a frozen teacher model; (ii) a translation module that maps student features into the teacher’s multimodal space; and (iii) a distillation objective combining local alignment with relational embedding regularization across regions. We validate MOCHA on personal object detection datasets, where the goal is to adapt a general detector to user-specific personal classes (e.g., distinguishing a generic dog from a particular user’s dog). MOCHA achieves significant improvements over state-of-the-art methods while retaining minimal inference cost, making it practical for resource-constrained deployments.
@@ 2 Related Work
Traditional knowledge distillation methods (Romero et al. 2015; Zhang et al. 2020; Passalis, Tzelepi, and Tefas 2020; Luo et al. 2019; Gou et al. 2024; Hao et al. 2022; Li et al. 2024; Tu et al. 2025) typically assume that teacher and student networks share similar architectures. The adoption of large transformer-based models (Vaswani et al. 2017; Dosovitskiy 2020; Devlin et al. 2019; Liu et al. 2021) has raised new challenges in transferring knowledge into efficient CNNs, which offer better trade-offs for deployment. Recent works (Liu et al. 2022; Hao et al. 2023) address this by projecting multi-scale features from teacher and student into a shared space for supervision, echoing shortcut-based strategies (Szegedy et al. 2014; Lee et al. 2014) to improve convergence stability.
These methods primarily aim to improve performance on the same task shared by teacher and student. In contrast, our approach aligns more closely with AuXFT (Barbato et al. 2024b), which leverages teacher features to guide a student on a related auxiliary task without degrading performance on the original one. Specifically, AuXFT addresses few-shot instance-level personalized object detection by injecting high-level teacher features into a lightweight detector to support prototype-based classification. Our method differs in three ways: (i) we leverage multimodal supervision from both visual and textual signals, unlike purely visual clues used in prior art; (ii) we distill compact region-level embeddings rather than dense intermediate activations, improving efficiency; and (iii) we explicitly regularize the embedding space to encourage geometric and relational alignment between individual instances.
@@
Several cross-modal learning strategies have been proposed to exploit complementary information across modalities (Xia et al. 2023; Barbato et al. 2024a) or to enforce consistency between predictions from different modalities (Jaritz et al. 2022, 2020). In our work, we transfer rich multimodal semantics from a vision-language model into a compact student detector to embed multimodal priors into a lightweight architecture.
@@
Personalization was first explored in Natural Language Processing (NLP) (Liu et al. 2023a; Vaswani et al. 2017; Devlin et al. 2019). Later, it was extended to computer vision, with early applications in personalized semantic segmentation (Zhang et al. 2021) and object detection (Michieli et al. 2024). The advent of foundation models such as CLIP (Radford et al. 2021) and SAM (Kirillov et al. 2023) has shifted personalization strategies toward prompt-based control of generalist architectures (Meng, Xia, and Ma 2024). For example, ViLD (Gu et al. 2022) aligns vision and language embeddings to enable open-vocabulary detection, paving the way for instance- and user-aware predictions through textual prompts. Similarly, SwissDINO (Paramonov et al. 2024), PerSAM (Zhang et al. 2024), Matcher (Liu et al. 2023b), and SegGPT (Wang et al. 2023) exploit prompt tuning and strong visual representations to tailor predictions to user preferences. Our work instead distills rich semantic features into a compact detector during server-side training, avoiding reliance on prompts or online adaptation and enabling efficient few-shot personalization for resource-constrained deployment.
@@ 3.1 Problem Setup
We begin with a standard detection model $\mathrm{m}_{S}(\cdot)=\mathrm{l}_{S}\circ\mathrm{g}_{S}$, where $\mathrm{l}_{S}(\cdot)$ is the detection head, $\mathrm{g}_{S}(\cdot)$ is the student backbone, and $\circ$ is the composition operator. We train it on a pretraining detection dataset using standard objectives $\mathcal{L}_{det}$ (Jocher, Chaurasia, and Qiu 2023). This stage results in a strong generic detector that identifies broad object categories but lacks personalized semantics or fine-grained embedding separation and serves as initialization for the student’s weights.
@@
Figure 3: Toy example. Effect of $\mathcal{L}_{\mathrm{emb}}$ on a set of ten 2D points optimized with respect to 3D ones. (a) 3D reference points, proxy for the teacher embeddings $u_{i}^{\prime}$. (b) Evolution of the 2D points (proxy for student embeddings $f^{\prime}_{A,i}$) updated via $\mathcal{L}_{\mathrm{emb}}$ from the 3D reference points $u_{i}^{\prime}$. $\star$ marks the original location, timesteps increase with color saturation. (c) Percent rate of 2D top-$k$ nearest neighbors (k-NN) that match those of the reference 3D distribution.
Few-Shot Personalization.
After feature distillation, the student backbone is frozen and used to extract multiscale region-level features $\{F_{j}\}_{j}=\mathrm{g}_{S}(X)$ from a personal detection dataset $\mathcal{D}_{f}$, labeled with bounding boxes $\mathcal{B}$ and class labels $\mathcal{C}_{f}$. The features are translated into the shared embedding space via the frozen translation module (Fig. 2) and fed to a prototype-based few-shot learner $\mathrm{p}(\cdot)$, like nearest class mean (Snell, Swersky, and Zemel 2017), which is trained and tested as in (Barbato et al. 2024b). More in detail, during personalization, a user provides a few labeled samples of personal instances of objects to the system (e.g., 1-5 samples). Later, during inference, the system automatically overrides the coarse predictions with the personal labels. This strategy enables the student to retain its compact architecture while adapting to novel visual concepts. Since no teacher inference or prompt engineering is required at personalization time, the method is efficient and suitable for deployment in real-world low-resource settings.
@@ 3.2 Multimodal Supervision Extraction
Without loss of generality, in our experiments, we instantiate the teacher with a LLaVa (Liu et al. 2023a) model, which includes a shared projection matrix $W_{T}$, in addition to $\mathrm{g}_{T}(\cdot)$ and $\mathrm{f}_{T}(\cdot)$. Each cropped image region $X_{i}$ is processed by the teacher’s visual encoder to obtain a visual embedding $Z_{V,i}$, projected into a joint text-vision representation space as $H_{V,i}=W_{T}Z_{V,i}$. Simultaneously, the class label $c_{i}$ is embedded as a textual query $Z_{Q,i}$ and projected via $H_{Q,i}=W_{T}Z_{Q,i}$. The pair $(H_{V,i},H_{Q,i})$ is passed to the frozen language model $\mathrm{f}_{T}$, which computes a sequence of tokens describing the interaction between the region and the class. We average the output along the temporal dimension to obtain a single semantic embedding $h_{i}\in\mathbb{R}^{d_{h}}$ representing the fused visual-linguistic description of the region $X_{i}$.
@@
Remarkably, the deviations closely follow a hyperbolic shape, allowing us to estimate their value based only on the channel index (detailed discussion in Appendix).
@@ 3.3 Student Detector and Feature Aggregation
To perform object-level supervision, we apply the same cropping strategy used for the teacher: for each ground-truth bounding box $b_{i}$, we first resize the feature maps to a common size to account for resolution differences, ensuring consistent spatial alignment. Then, we extract the corresponding region-aligned features $F_{A,j,i}$ from each resolution level $j$, and apply spatial average pooling to obtain fixed-size descriptors $f_{A,j,i}\in\mathbb{R}^{d_{s}}$.
@@
Without loss of generalization, we adopt YOLOv8 (Jocher, Chaurasia, and Qiu 2023) as the student model pretrained on the COCO dataset (Lin et al. 2014) and then fine-tuned in MOCHA.
@@ 3.4 Training Objectives and Strategy
The primary objective enforces direct alignment between translated student features $f^{\prime}_{A,i}$ and their corresponding teacher vector $u^{\prime}_{i}$. We define the distillation loss as the average of $\ell_{1}$ and $\ell_{2}$ distances, as in (Barbato et al. 2024b):
@@
Beyond individual alignment, we encourage the student to preserve the global geometry of the teacher’s embedding space. To this end, we define a relational embedding loss based on pairwise Euclidean distances between all student and teacher features.
Given the translated student features $\{f^{\prime}_{A,i}\}_{i}$ and the corresponding teacher targets $\{u^{\prime}_{i}\}_{i}$, we compute two pairwise distance matrices $D_{f\!f}$ and $D_{uu}$. We convert them into normalized distributions using temperature-scaled softmax:
$P_{f\!f}=\mathrm{softmax}\left(-\frac{D_{f\!f}}{\tau}\right),\ \ \ P_{uu}=\mathrm{softmax}\left(-\frac{D_{uu}}{\tau}\right),$ (7)
where $\tau$ is a hyperparameter controlling distribution sharpness. The embedding loss is then defined as the cross-entropy between the two distributions:
@@
where the $\lambda$ coefficients regulate the contribution of each auxiliary term. By integrating multimodal feature supervision with structural regularization, MOCHA enables robust semantic transfer into compact detectors, improving performance in low-resource and personalized scenarios.
Figure 4: PCA study. Average mAP50-95 across all evaluation datasets varying the feature dimension $d_{t}$.
@@ 4 Experiments
In this section, we introduce the experimental setup to test MOCHA’s ability to effectively distill knowledge from a multimodal teacher into a lightweight student model, with the goal of achieving robust performance in few-shot personalized object detection across four personal datasets. We begin by discussing general analyses, followed by comparisons with state-of-the-art approaches, and conclude by presenting ablation studies.
@@ 4.1 Experimental Setup
For pretraining, we adopt COCO (Lin et al. 2014) and OpenImagesV7 (Kuznetsova et al. 2020). Also, we consider starting from AuXFT pretrained weights (which start from COCO and fine-tune on OpenImages). Distillation is performed on OpenImages. Considered personalization datasets are: PerSeg (Zhang et al. 2024), POD (Barbato et al. 2024b), CORe50 (Lomonaco and Maltoni 2017), and iCubWorld (Fanello et al. 2013).
@@
Our approach uses YOLOv8n as the student model and LLaVa-1.5-7B as the teacher, which includes CLIP with ViT-B/32 as the visual encoder and LLaMA 7B as the text encoder. For comparison, we also include: (i) the text encoder from LLaVa, i.e., LLaMA 7B (ii) the visual encoder from LLaVa, i.e., CLIP, and (iii) DINOv2 (Oquab et al. 2023). Moreover, we compare against other distillation-based approaches: OFA (Hao et al. 2023), knowledge distillation at the output space via KL divergence (KL) (Touvron et al. 2021) and its combination with embedding space distillation via MSE (MSE + KL) (Romero et al. 2015), open-vocabulary detectors (VILD) (Gu et al. 2022), and the most relevant personalization method to date (AuXFT).
@@
Figure 5: Feature similarity between the $F_{A}$ embeddings and the teacher target $u^{\prime}_{i}$ is encoded in the R/G/B channels, one for each object in the input scene from POD dataset.
@@
YOLO 69.6 $\pm 3.4$ 35.4 $\pm 3.4$ 41.3 $\pm 0.0$ 31.5 $\pm 5.5$ 42.8 $\pm 7.8$ 33.3 $\pm 2.7$ 48.6 $\pm 2.4$ 43.2
CLIP ($z_{V,i}$) YOLO 75.5 $\pm 3.5$ 40.9 $\pm 3.2$ 51.7 $\pm 0.0$ 35.0 $\pm 3.9$ 47.3 $\pm 5.7$ 40.6 $\pm 2.9$ 60.0 $\pm 2.3$ 50.7
LLaVa ($h_{i}$) YOLO 73.0 $\pm 3.6$ 37.1 $\pm 4.2$ 47.6 $\pm 0.0$ 33.9 $\pm 3.7$ 45.9 $\pm 5.4$ 37.4 $\pm 3.1$ 55.8 $\pm 2.6$ 48.0
LLaVa ($h_{i}$) + CLIP ($z_{V,i}$) YOLO 82.0 $\pm 3.4$ 45.7 $\pm 3.5$ 58.1 $\pm 0.0$ 34.5 $\pm 4.1$ 47.8 $\pm 5.9$ 42.9 $\pm 3.1$ 62.7 $\pm 2.2$ 54.0
Table 1: Oracle results using ground truth bounding boxes on personal datasets. Top block: performance when using frozen teacher embeddings directly in the personalization stage. Bottom block: performance of YOLO distilled with the teacher signals.
@@
YOLO 41.2 $\pm 2.6$ 23.6 $\pm 2.4$ 30.4 $\pm 0.0$ 57.8 $\pm 6.3$ 67.3 $\pm 7.7$ 51.2 $\pm 3.0$ 68.4 $\pm 2.4$ 48.6
DINO AuXFT YOLO 48.8 $\pm 3.4$ 31.5 $\pm 2.7$ 38.8 $\pm 0.0$ 58.8 $\pm 6.4$ 69.3 $\pm 6.0$ 55.0 $\pm 3.0$ 74.5 $\pm 2.5$ 53.8
CLIP ($z_{V,i}$) ViLD YOLO 44.4 $\pm 3.2$ 24.8 $\pm 2.8$ 33.3 $\pm 0.0$ 54.6 $\pm 5.2$ 64.9 $\pm 4.9$ 57.7 $\pm 3.0$ 70.4 $\pm 2.4$ 50.9
LLaVa ($u_{i}$) KL div. YOLO 43.7 $\pm 2.9$ 23.4 $\pm 2.5$ 27.8 $\pm 0.0$ 54.4 $\pm 5.0$ 65.2 $\pm 5.0$ 54.7 $\pm 3.0$ 69.5 $\pm 2.3$ 49.0
LLaVa ($u_{i}$) KL div. + MSE YOLO 50.1 $\pm 3.0$ 27.6 $\pm 2.4$ 30.5 $\pm 0.0$ 53.3 $\pm 4.5$ 65.7 $\pm 4.6$ 61.1 $\pm 2.9$ 78.5 $\pm 1.9$ 52.1
LLaVa ($u_{i}$) OFA YOLO 47.2 $\pm 3.4$ 23.3 $\pm 2.6$ 28.0 $\pm 0.0$ 55.3 $\pm 4.3$ 63.6 $\pm 4.3$ 60.4 $\pm 2.5$ 72.5 $\pm 2.4$ 50.1
CLIP ($z_{V,i}$) MOCHA (visual) YOLO 53.0 $\pm 2.8$ 26.7 $\pm 3.1$ 37.5 $\pm 0.0$ 57.2 $\pm 4.7$ 67.2 $\pm 4.8$ 62.6 $\pm 2.8$ 76.1 $\pm 2.3$ 54.6
LLaVa ($h_{i}$) MOCHA (text) YOLO 52.3 $\pm 2.8$ 28.8 $\pm 2.6$ 33.7 $\pm 0.0$ 56.2 $\pm 4.7$ 67.1 $\pm 4.7$ 62.8 $\pm 2.8$ 76.8 $\pm 2.0$ 54.0
LLaVa ($u_{i}$) MOCHA (COCO) YOLO 55.4 $\pm 3.3$ 33.9 $\pm 2.8$ 38.9 $\pm 0.0$ 55.8 $\pm 4.7$ 65.4 $\pm 4.7$ 60.6 $\pm 3.3$ 77.6 $\pm 2.5$ 56.0
LLaVa ($u_{i}$) MOCHA (AuXFT) YOLO 59.1 $\pm 3.3$ 36.3 $\pm 3.3$ 45.9 $\pm 0.0$ 60.9 $\pm 4.4$ 70.6 $\pm 4.5$ 61.4 $\pm 3.1$ 77.0 $\pm 2.3$ 58.7
Table 2: Few-shot personalized detection results on personal datasets. First row: YOLO baseline with no distillation. First block: MOCHA compared against prior state-of-the-art supervision methods. Second block: evaluation of different MOCHA variants.

### v4

@@ Abstract
Personalized object detection aims to adapt a general-purpose detector to recognize user-specific instances from only a few examples. Lightweight models often struggle in this setting due to their weak semantic priors, while large vision-language models (VLMs) offer strong object-level understanding but are too computationally demanding for real-time or on-device applications. We introduce MOCHA (Multi-modal Objects-aware Cross-arcHitecture Alignment), a distillation framework that transfers multimodal region-level knowledge from a frozen VLM teacher into a lightweight vision-only detector. MOCHA extracts fused visual and textual teacher’s embeddings and uses them to guide student training through a dual-objective loss that enforces accurate local alignment and global relational consistency across regions. This process enables efficient transfer of semantics without the need for teacher modifications or textual input at inference. MOCHA consistently outperforms prior baselines across four personalized detection benchmarks under strict few-shot regimes, yielding a +10.1 average improvement, with minimal inference cost.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) such as CLIP [36], Flamingo [1], and LLaVa [23] have demonstrated remarkable zero-shot and open-vocabulary capabilities, owing to their rich representations and scale. These models, however, are typically large and computationally intensive, limiting their applicability in real-time or resource-constrained scenarios such as mobile devices (e.g., smartphones or robots). In contrast, lightweight architectures such as YOLO [17, 16] offer fast and memory-efficient object detection, at the cost of degraded performance in low-data regimes and vulnerability to neural collapse [19, 33], where features become indistinguishably aligned and lose their discriminative power. In this work, we aim to bridge this gap by proposing MOCHA, a knowledge distillation approach that transfers object-centric multimodal embeddings from a vision-language teacher into a compact vision-only student detector. MOCHA builds on the observation that semantically related concepts tend to exhibit similar structures in the embedding space across modalities under a well-generalized backbone [13, 8, 4]. By aligning and regularizing the student’s feature space, it achieves improved generalization in few-shot detection settings. Our approach consists of three stages (Fig. 1): (1) pretraining a student detector, (2) distilling rich joint visual-textual features from a frozen teacher, and (3) performing few-shot personalization with a frozen student and a prototype-based classifier [38].
@@
Our primary contribution lies in the distillation stage, where we introduce: (i) extraction of joint visual-language features from a frozen teacher, using both image regions and class labels to provide strong object-level supervision; (ii) a translation module that maps the student’s visual features into the teacher’s multimodal space; and (iii) a distillation objective that combines local feature alignment with a relational constraint, helping the student match the teacher’s semantics and get relationships between object regions. We validate MOCHA on personal object detection datasets, where the goal is to adapt a general detector to user-specific personal classes (e.g., distinguishing a generic dog from a particular user’s dog). MOCHA achieves significant improvements over state-of-the-art methods while retaining minimal inference cost, making it practical for resource-constrained deployment.
@@ 2 Related Work
Traditional knowledge distillation methods [37, 45, 35, 28, 9, 11, 21, 40] typically assume that teacher and student networks share similar architectures. The adoption of large transformer-based models [41, 6, 5, 26] has raised new challenges in transferring knowledge into efficient CNNs, which offer better trade-offs for deployment. Recent works [24, 12] address this by projecting multi-scale features from teacher and student into a shared space for supervision. These methods primarily aim to improve performance on the same task shared by teacher and student. In contrast, our approach aligns more closely with AuXFT [3], which leverages teacher features to guide a student on a related auxiliary task without degrading performance on the original one. Specifically, AuXFT addresses few-shot instance-level personalized object detection by injecting high-level teacher features into a lightweight detector to support prototype-based classification. To the best of our knowledge, AuXFT is the only prior work exploring a similar setup. Our method differs in three ways: (i) we leverage multimodal supervision from both visual and textual signals, unlike purely visual clues used in prior art; (ii) we distill compact region-level embeddings rather than dense intermediate activations, improving efficiency; and (iii) we explicitly regularize the embedding space to encourage geometric and relational alignment between individual instances.
@@
Several cross-modal learning strategies have been proposed to exploit complementary information across modalities [44, 2] or to enforce consistency between predictions from different modalities [15, 14]. In our work, we transfer rich multimodal semantics from a vision-language model into a compact student detector to embed multimodal priors into a lightweight architecture.
@@
Personalization was first explored in Natural Language Processing (NLP) [23, 41, 5]. Later, it was extended to computer vision, with early applications in personalized semantic segmentation [47] and object detection [31]. The advent of foundation models such as CLIP [36] and SAM [18] has shifted personalization strategies toward prompt-based control of generalist architectures [30]. Building on this trend, several methods leverage strong visual or multimodal representations to guide instance-aware predictions through textual or visual prompts, including open-vocabulary detectors like ViLD [10] and prompt-driven approaches such as SwissDINO [34], PerSAM [46], Matcher [25], and SegGPT [42]. While these approaches allow for zero-shot recognition, they are computationally demanding and do not directly address the few-shot instance-level recognition scenario considered here. Our work instead distills rich semantic features into a compact detector during server-side training, avoiding reliance on textual prompts at test time (that would add significant computational costs) or online adaptation, and enabling efficient few-shot personalization for resource-constrained deployment.
@@ 3.1 Problem Setup
We begin with a standard detection model $\mathrm{m}_{S}(\cdot)=\mathrm{l}_{S}\circ\mathrm{g}_{S}$, where $\mathrm{l}_{S}(\cdot)$ is the detection head, $\mathrm{g}_{S}(\cdot)$ is the student backbone, and $\circ$ is the composition operator. We train it on a pretraining detection dataset using standard objectives $\mathcal{L}_{det}$ [17]. This stage results in a strong generic detector that identifies broad object categories but lacks personalized semantics or fine-grained embedding separation and serves as initialization for the student’s weights.
@@
Few-Shot Personalization.
After feature distillation, the student backbone is frozen and used to extract multiscale region-level features $\{F_{j}\}_{j}=\mathrm{g}_{S}(X)$ from a personal detection dataset $\mathcal{D}_{f}$, labeled with bounding boxes $\mathcal{B}$ and class labels $\mathcal{C}_{f}$. The features are translated into the shared embedding space via the frozen translation module (Fig. 2) and fed to a prototype-based few-shot learner $\mathrm{p}(\cdot)$, like nearest class mean [38], which is trained and tested as in [3]. More in detail, during personalization, a user provides a few labeled samples of personal instances of objects to the system (e.g., 1-5 samples). Later, during inference, the system automatically overrides the coarse predictions with the personal labels. This strategy enables the student to retain its compact architecture while adapting to novel visual concepts. Since no teacher inference or prompt engineering is required at personalization time, the method is efficient and suitable for deployment in real-world low-resource settings.
@@
Figure 3: Effect of $\mathcal{L}_{\mathrm{emb}}$ on a set of ten 2D points optimized with respect to 3D ones. (a) 3D reference points, proxy for the teacher embeddings $u_{i}^{\prime}$. (b) Evolution of the 2D points (proxy for student embeddings $f^{\prime}_{A,i}$) updated via $\mathcal{L}_{\mathrm{emb}}$ from the 3D reference points $u_{i}^{\prime}$. $\star$ marks the original location, timesteps increase with color saturation. (c) Percent rate of 2D top-$k$ nearest neighbors (k-NN) that match those of the reference 3D distribution.
@@ 3.2 Multimodal Supervision Extraction
Without loss of generality, in our experiments, we instantiate the teacher with a LLaVa [23] model, which includes a shared projection matrix $W_{T}$, in addition to $\mathrm{g}_{T}(\cdot)$ and $\mathrm{f}_{T}(\cdot)$. Each cropped image region $X_{i}$ is processed by the teacher’s visual encoder to obtain a visual embedding $Z_{V,i}$, projected into a joint text-vision representation space as $H_{V,i}=W_{T}Z_{V,i}$. Simultaneously, the class label $c_{i}$ is embedded as a textual query $Z_{Q,i}$ and projected via $H_{Q,i}=W_{T}Z_{Q,i}$. The pair $(H_{V,i},H_{Q,i})$ is passed to the frozen language model $\mathrm{f}_{T}$, which computes a sequence of tokens describing the interaction between the region and the class. We average the output along the temporal dimension to obtain a single semantic embedding $h_{i}\in\mathbb{R}^{d_{h}}$ representing the fused visual-linguistic description of the region $X_{i}$.
@@
Remarkably, the deviations closely follow a hyperbolic shape, allowing us to estimate their value based only on the channel index (detailed discussion in the Appendix).
@@ 3.3 Student Detector and Feature Aggregation
To perform object-level supervision, we apply the same cropping strategy used for the teacher: for each ground truth bounding box $b_{i}$, we first resize the feature maps to a common size to account for resolution differences, ensuring consistent spatial alignment. Then, we extract the corresponding region-aligned features $F_{A,j,i}$ from each resolution level $j$, and apply spatial average pooling to obtain fixed-size descriptors $f_{A,j,i}\in\mathbb{R}^{d_{s}}$.
@@ 3.4 Training Objectives and Strategy
The primary objective enforces direct alignment between translated student features $f^{\prime}_{A,i}$ and their corresponding teacher vector $u^{\prime}_{i}$. We define the distillation loss as the average of $\ell_{1}$ and $\ell_{2}$ distances, as in [3]:
@@
Beyond individual alignment, we encourage the student to preserve the global geometry of the teacher’s embedding space. To this end, we define a relational embedding loss based on pairwise Euclidean distances between all student and teacher features. Given the translated student features $\{f^{\prime}_{A,i}\}_{i}$ and the corresponding teacher targets $\{u^{\prime}_{i}\}_{i}$, we compute pairwise distance matrices $D_{f\!f},D_{uu}$ using normalized Euclidean distances. We remove self-distances by discarding diagonal entries and convert into probability distributions using softmax:
$P_{f\!f}=\mathrm{softmax}\left(-D_{f\!f}\right),\ \ P_{uu}=\mathrm{softmax}\left(-D_{uu}\right).$ (7)
The embedding loss is then defined as the cross-entropy between the two distributions:
@@
where the $\lambda$ coefficients regulate the contribution of each auxiliary term. This combined objective ensures that the student learns both task-relevant representations and the semantic structure transferred from the teacher.
Figure 4: mAP50-95 at different PCA dimensionality. Average score across all evaluation datasets varying feature dimension $d_{t}$.
@@ 4 Experiments
We evaluate MOCHA in the context of few-shot personalized object detection, assessing its ability to transfer multimodal knowledge from a large VLM teacher into lightweight student detectors. Our experiments cover four personal datasets and analyze performance from three complementary perspectives: general behavior and design insights, comparison with state-of-the-art methods, and detailed ablation studies. Results are reported after introducing the experimental setup and presenting preliminary analyses.
@@ 4.1 Experimental Setup
For pretraining, we adopt COCO [22] and OpenImagesV7 [20]. Also, we consider starting from AuXFT pretrained weights (which in turn start from COCO and fine-tune on OpenImages). Distillation is performed on OpenImages. Considered personalization datasets are: PerSeg [46], POD [3], CORe50 [27], and iCubWorld [7].
@@
Our approach uses YOLOv8n ($\approx 3.2$M params) as the student model and LLaVa-1.5-7B ($\approx 7.2$B params) as the teacher, which includes CLIP with ViT-B/32 as the visual encoder and LLaMA 7B as the text encoder. For comparison, we also include: (i) the text encoder from LLaVa, i.e., LLaMA 7B ($\approx 7$B params) (ii) the visual encoder from LLaVa, i.e., CLIP ($\approx 0.2$B params), and (iii) DINOv2 [32] ($\approx 22$M params). Moreover, we compare against other distillation-based approaches: OFA [12], knowledge distillation at the output space via KL divergence (KL) [39] and its combination with embedding space distillation via MSE (MSE + KL) [37], open-vocabulary detectors (VILD) [10], and the most relevant personalization method to date (AuXFT). Finally, we perform ablations on the student model including YOLOv11n [16] ($\approx 2.6$M params) and RT-DETR-l [29] ($\approx 45$M params).
@@
Figure 5: Feature similarity between the $F_{A}$ embeddings and the teacher target $u^{\prime}_{i}$ encoded in the R/G/B channels, one for each object in the input scene from POD dataset.
@@
YOLOv8n 69.6 $\pm 3.4$ 35.4 $\pm 3.4$ 41.3 $\pm 0.0$ 31.5 $\pm 5.5$ 42.8 $\pm 7.8$ 33.3 $\pm 2.7$ 48.6 $\pm 2.4$ 43.2
CLIP ($z_{V,i}$) YOLOv8n 75.5 $\pm 3.5$ 40.9 $\pm 3.2$ 51.7 $\pm 0.0$ 35.0 $\pm 3.9$ 47.3 $\pm 5.7$ 40.6 $\pm 2.9$ 60.0 $\pm 2.3$ 50.7
LLaVa ($h_{i}$) YOLOv8n 73.0 $\pm 3.6$ 37.1 $\pm 4.2$ 47.6 $\pm 0.0$ 33.9 $\pm 3.7$ 45.9 $\pm 5.4$ 37.4 $\pm 3.1$ 55.8 $\pm 2.6$ 48.0
LLaVa ($h_{i}$) + CLIP ($z_{V,i}$) YOLOv8n 82.0 $\pm 3.4$ 45.7 $\pm 3.5$ 58.1 $\pm 0.0$ 34.5 $\pm 4.1$ 47.8 $\pm 5.9$ 42.9 $\pm 3.1$ 62.7 $\pm 2.2$ 54.0
Table 1: Oracle results using ground truth bounding box coordinates on personal datasets. Top block: performance when using frozen teacher embeddings directly in the personalization stage. Bottom block: performance of student distilled with the teacher signals.
@@
YOLOv8n 41.2 $\pm 2.6$ 23.6 $\pm 2.4$ 30.4 $\pm 0.0$ 57.8 $\pm 6.3$ 67.3 $\pm 7.7$ 51.2 $\pm 3.0$ 68.4 $\pm 2.4$ 48.6
DINO AuXFT YOLOv8n 48.8 $\pm 3.4$ 31.5 $\pm 2.7$ 38.8 $\pm 0.0$ 58.8 $\pm 6.4$ 69.3 $\pm 6.0$ 55.0 $\pm 3.0$ 74.5 $\pm 2.5$ 53.8
CLIP ($z_{V,i}$) ViLD YOLOv8n 44.4 $\pm 3.2$ 24.8 $\pm 2.8$ 33.3 $\pm 0.0$ 54.6 $\pm 5.2$ 64.9 $\pm 4.9$ 57.7 $\pm 3.0$ 70.4 $\pm 2.4$ 50.9
LLaVa ($u_{i}$) KL div. YOLOv8n 43.7 $\pm 2.9$ 23.4 $\pm 2.5$ 27.8 $\pm 0.0$ 54.4 $\pm 5.0$ 65.2 $\pm 5.0$ 54.7 $\pm 3.0$ 69.5 $\pm 2.3$ 49.0
LLaVa ($u_{i}$) KL div. + MSE YOLOv8n 50.1 $\pm 3.0$ 27.6 $\pm 2.4$ 30.5 $\pm 0.0$ 53.3 $\pm 4.5$ 65.7 $\pm 4.6$ 61.1 $\pm 2.9$ 78.5 $\pm 1.9$ 52.1
LLaVa ($u_{i}$) OFA YOLOv8n 47.2 $\pm 3.4$ 23.3 $\pm 2.6$ 28.0 $\pm 0.0$ 55.3 $\pm 4.3$ 63.6 $\pm 4.3$ 60.4 $\pm 2.5$ 72.5 $\pm 2.4$ 50.1
CLIP ($z_{V,i}$) MOCHA (visual) YOLOv8n 53.0 $\pm 2.8$ 26.7 $\pm 3.1$ 37.5 $\pm 0.0$ 57.2 $\pm 4.7$ 67.2 $\pm 4.8$ 62.6 $\pm 2.8$ 76.1 $\pm 2.3$ 54.6
LLaVa ($h_{i}$) MOCHA (text) YOLOv8n 52.3 $\pm 2.8$ 28.8 $\pm 2.6$ 33.7 $\pm 0.0$ 56.2 $\pm 4.7$ 67.1 $\pm 4.7$ 62.8 $\pm 2.8$ 76.8 $\pm 2.0$ 54.0
LLaVa ($u_{i}$) MOCHA (COCO) YOLOv8n 55.4 $\pm 3.3$ 33.9 $\pm 2.8$ 38.9 $\pm 0.0$ 55.8 $\pm 4.7$ 65.4 $\pm 4.7$ 60.6 $\pm 3.3$ 77.6 $\pm 2.5$ 56.0
LLaVa ($u_{i}$) MOCHA (AuXFT) YOLOv8n 59.1 $\pm 3.3$ 36.3 $\pm 3.3$ 45.9 $\pm 0.0$ 60.9 $\pm 4.4$ 70.6 $\pm 4.5$ 61.4 $\pm 3.1$ 77.0 $\pm 2.3$ 58.7
Table 2: Few-shot personalized detection results on personal datasets. First row: student baseline with no distillation. First block: MOCHA compared against prior state-of-the-art supervision methods. Second block: evaluation of different MOCHA variants.